Consider a concept learning example of 5 attributes. Attributes can take one of (5,3,2,2,6) values respectively. Determine the sizes of Instance Space, Hypothesis Space ( used by Candidate-Elimination Algorithm) and the Space of all possible Concepts.
Answer #1

All the attributes has 4 distinct possible values i.e 5,3,2 and 6.

Therefore,

Instance Space = 4*4*4*4*4 = 1024.

Hypothesis space :

In this we add to each possible value a [null] so, now number of possible values of each attribute is 5.

= 1 + 5*5*5*5*5 = 3125 + 1 = 3126.

Space of all possible concepts:

In this we add to each possible value a [null] and a [0] so, now number of possible values of each attribute is 6

= 6*6*6*6*6 = 7776.